OpenStudy (anonymous):

Andrew estimated the weight of his dog to be 60 lb. The dog’s actual weight was 68 lb. What was the percent error in Andrew’s estimate? Round your answer to the nearest tenth of a percent.

OpenStudy (darthvader2900):

Is this for a test? If it is, can I have the unit and lesson number?

OpenStudy (anonymous):

@563blackghost can you help me on this one please

563blackghost (563blackghost):

\[\frac{ (60-68) }{ 68 }\times 100percent\]

563blackghost (563blackghost):

https://www.mathsisfun.com/numbers/percentage-error.html

OpenStudy (anonymous):

im confused? @563blackghost

563blackghost (563blackghost):

Ok...lets work through it together... I setted up the equation so all we need to do is simplify it... sooo... we first will subtract 60-68 what is the answer?

OpenStudy (anonymous):

-8

563blackghost (563blackghost):

Correct! We will ignore the negative sign since we are trying to find a percent...now we have to divide 8 by 63... \[\frac{ 8 }{ 63 }\] so whats the answer?

OpenStudy (anonymous):

0.126

563blackghost (563blackghost):

sorry! I meant 8 divided by 68...\[\frac{ 8 }{ 68 }\] plz forgive me...

OpenStudy (anonymous):

lol its ok 0.117

563blackghost (563blackghost):

Correct! Now we need to times .117 by 100% but the easiest way to do that is to move the decimal sign to the right 2 times... so whats the percent?

OpenStudy (anonymous):

11.7%?

563blackghost (563blackghost):

Correct!!! the percent error is 11.7% !

OpenStudy (anonymous):

I'm not trying to get you to answer these for me but few more if that's ok its honestly helping a lot any of them I can get right I hope I can pass this I cant retake it

563blackghost (563blackghost):

Sure I can help with more!

OpenStudy (anonymous):

What is the exact area of a circle whose diameter is 24 cm?

563blackghost (563blackghost):

To find the area of a circle you use this formula...\[A=\pi r^2\] We know the diameter but what can we do to the diameter to get the radius?

OpenStudy (anonymous):

divide by 2 and get 12?

563blackghost (563blackghost):

Correct! Now we know the radius so lets input it into the formula... \[A=3.14(12^2)\] we substitute pi with 3.14.....We do order of operations so what is...\[12^2=?\]

OpenStudy (anonymous):

144

563blackghost (563blackghost):

Yup! so...\[A=3.14(144)\] Now all we need to do is multiply 3.14 by 144 which will give us the area of the circle....

OpenStudy (anonymous):

452.16

563blackghost (563blackghost):

Correct! So the area of the circle is...\[452.16cm^2\]

OpenStudy (anonymous):

that isnt one of our choices though? :/

OpenStudy (anonymous):

its 576 144 48 or 24

563blackghost (563blackghost):

hmm...we did the process correct... O.o

OpenStudy (anonymous):

maybe its 144

OpenStudy (anonymous):

maybe we didnt have to multiply the 3.14 again so it would stay 144 O.o

563blackghost (563blackghost):

probably...but the question says " exact value of the area"....its ur choice

OpenStudy (anonymous):

sorry im back are you still on? @563blackghost